2m+4-3m=8(m-1) m= ?
Answer by ReadingBoosters(3246)

Starting with
2m + 4 - 3m = 8(m-1)m = ?
What is the sign between the )m?
Distribute the 8
2m + 4 - 3m = 8m -8 ? m
Combine like terms
-m + 4 = 8m?m - 8
Isolate m
4 + 8 = 8m?m + m
12 = 8m?m + m
Divide both sides by the m coefficient (the number in front of m) to solve for m
Repost if you'd like to check your answer.
........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.